Assume f is a one-to-one function. If f(x)=3-6xf , find f^-1(33)

Answer by ikleyn(52781)

.
Assume f is a one-to-one function. If f(x)=3-6xf , find f^-1(33)
~~~~~~~~~~~~~~~~~~


The problem is presented very unprofessionally in this post.

If f(x) = 3 - 6x, then f(x) is a linear function of x. Since a coefficient at x is not equal to zero,
this function is one-to-one without any assumption: you don't need make this assumption - it is TRUE without any assumption.

So, the correct formulation should be in this form:

        If f(x)=3-6x, find f^-1(33).

In Math, when you formulate a problem, all the words / (the terms) should be used correctly and must be placed in a right order.

Also, unnecessary words should not obscure a meaning of the problem and should not interfere with understanding.

It is the same as in music (classic music, I mean . . . )


///////////////


OK, now I will solve the problem. So, the question is

    +----------------------------------+
    |    If f(x)=3-6x, find f^-1(33).  |
    +----------------------------------+


To get the answer, you do not need express f^(-1) (x) explicitly as a function of x.


In this problem, the KEY POINT is to understand that the question is equivalent to finding x from equation

    f(x) = 33,  which is  3 - 6x = 33.


From this equation, you have

    3 - 33 = 6x

      -30  = 6x

       x   = (-30)/6 = -5.


ANSWER.  f^(-1)(33) = -5.


CHECK.   f(-5) = 3 - 6(-5) = 3 + 30 = 33.   ! Precisely correct !

Solved, answered and explained, including all incoming details.


////////////

The lesson to learn from my post


        (1)   To get the answer,  you do not need express  f^(-1) (x)  explicitly as a function of  x.

        (2)   In this problem, the question is equivalent to finding x from equation for a base function.